OpenStudy (anonymous):

evaluate: (i=complex number) (-3+2i)/8i

OpenStudy (ash2326):

@a.galvan2 We know that \[i^2=-1\] or \[\sqrt {-1}=i\] we have \[\frac{-3+2i}{8i}\] Multiply and divide by i \[\frac{-3+2i}{8i}\times \frac i i\] now we get \[\frac{-3i+2i^2}{8i^2}\] Do you get till here?

OpenStudy (anonymous):

so... 3i+2/8?

OpenStudy (ash2326):

yeah, you're correct Good work:D

OpenStudy (anonymous):

thank you! I think that's what i did, but was sure it wasn't that easy...

OpenStudy (ash2326):

Yeah, it happens sometimes:)